OpenStudy (anonymous):

help really stressed A juggler is performing her act using several balls. She throws the balls up at an initial height of 4 feet, with a speed of 15 feet per second. If the juggler doesn't catch one of the balls, about how long does it take the ball to hit the floor? Hint: Use H(t) = −16t2 + vt + s.

OpenStudy (anonymous):

please help me

Miracrown (miracrown):

\[h(t) = -16t^2 + vt + S\] We have an initial height of 4 feet in the equation, this is s ..we are also given a speed of 15 feet/second

Mehek (mehek14):

I think the quadratic function will work for this

Mehek (mehek14):

after plugging in everything

Miracrown (miracrown):

in the equation, this is V and we are interested in the amount of time it iwll take for the ball to drop tot he ground at this time, height will be zero \[h(t)=−16t^2+vt+S \space = 0\]

Miracrown (miracrown):

so what we have here is a quadratic equation, we can solve this using the quadratic formula like @Mehek14 said

Mehek (mehek14):

\(\bf{x=\dfrac{-b±\sqrt{b^2-4ac}}{2a}}\)

Mehek (mehek14):

original formula is \(\bf{f(x)=ax^2+bx+c}\)

Mehek (mehek14):

for a, you have -16 for b, you have 15 and for c, you have 4 @jaredhm29 can you plug that into the quadratic formula?

Mehek (mehek14):

@jaredhm29 are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ive had a stuggle in this segment

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

can you plug a, b , and c into the quadratic formula?

OpenStudy (anonymous):

idk what u mean sry

Mehek (mehek14):

\(\bf{x=\dfrac{-b±\sqrt{b^2-4ac}}{2a}}\)

Mehek (mehek14):

for a, you have -16 for b, you have 15 and for c, you have 4

Mehek (mehek14):

\(\bf{x=\dfrac{-15±\sqrt{15^2-4*(-16)*4}}{2*(-16)}}\)

Mehek (mehek14):

first thing what is \(15^2=?\)

Mehek (mehek14):

@jaredhm29 ?

OpenStudy (anonymous):

@Theloshua

OpenStudy (anonymous):

please help me @Miracrown

OpenStudy (anonymous):

@zzr0ck3r @Miracrown

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

As @Miracrown said, \(\large h(t) = -16t^2 + vt + S\) is a standard kinematics equation giving the height h above datum (ground) at time t (therefore h(t)) as a function of t. v is the initial velocity (up = positive)=15 ft/s, and S= initial location (4 ft above datum= ground). -16 is actually acceleration due to gravity (32.2 ft/s^2) divided by 2, negative because acceleration is downwards. Substitute these values, you will get the same equation as @mehek14 showed you. Solving for t and reject the negative root gives you the time required.